40 mL fog water sample collected. Marine layer returned as verified by NOAA NWS
40 mL rain water sample collected. Tropical moisture and rain verified by NOAA NWS
Arboretum and Twin Gates were downloaded on 9/6/17. Data was only being logged until 8/23
Younger Lagoon data was collected 9/7/17, and was still actively logging at the time of collection.
Big fog event. Collector came on around midnight and about 300 mL was collected.
Collector came on around 7 am but was still able to collect about 25 mL of fog water
Tropical storm on Monday and it rained. The collector came on at around 11:30 am and went off at about 1:30 pm. No sample was collected.
The fog collector came on for 3 minutes this morning at 8:43 am. No samples was collected. The collector was clean with soap and water and dried. No blank was taken.
(what was going on for 3 minutes on the roof?)
The fog collector came on at 12:48 am last night and turned off at 10:53 am this morning. 180mL of sample was collected. The collector was washed with soap and water and a blank was taken. Brian set up another light incubation experiment. Hopefully, non of the jars break this time!
The fog collector turned on at ~12am and turned off at 9:58 am. ~10mL of sample was collected. The collector was cleaned with soap and water and a blank was taken.
We noticed that the fog collector optical sensor was having trouble turning on, so we put Anti-Fog liquid on the metal trough to see if the water collected from the passive collector would roll off better.
It was very overcast here at UCSC so we didn’t get a chance to see the eclipse. The collector was been on since 6:00 am the of the previous day. It finally turn off at 2:00pm today. About 20mL of sample was collected. It was washed with soap and water, then a blank was taken.